Circular No. 5838

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J1008–57 J. E. Grove, R. L. Kroeger, B. Phlips, J. D. Kurfess, and W. N. John-

son, for the OSSE team; M. H. Finger, R. B. Wilson, B. A. Harmon, M. T. Stollberg, B. C. Rubin, and G. J. Fishman, for the BATSE team; and N. Gehrels, for the Compton Gamma-Ray Observatory, report: “An improved position for the recently discovered x-ray pulsar GRO J1008–57 (IAUC 5836) has been obtained by using OSSE scans of the region along with a re-evaluation of the BATSE Earth occultation times. The 90-percent confidence region is bounded by a box with the following corners: α = 10^{h}03^{m}, δ = −58^{o}25′ (equinox 2000.0); 10^{h}03^{m}, −58^{o}50′; 10^{h}13^{m}, −58^{o}20′; 10^{h}13^{m}, −57^{o}55′. The source flux observed by BATSE in the energy range 20-120 keV has remained within 15 percent of the value reported on IAUC 5836 from July 19 to 28.”

NOVA AQUILAE 1993 A. Evans and T. Naylor, Keele University; and L. Koch-Miramond,

Centre d’Etudes de Saclay, report: “We made polarimetric observations of N Aql 1993, which J. Elias et al. (IAUC 5831) suggest has entered a dust phase. The data, which were obtained in white light (roughly covering the region of the UBV filters) with the 2-m Bernard Lyot telescope at Pic du Midi Observatory, yield the following values: July 19.0 UT, P = 1.75 ± 0.15 percent, p.a. 42^{o} ± 2^{o}; July 23.9, 1.90 ± 0.15 percent, 48^{o} ± 8^{o}. The high degree of polarization suggests an intrinsic component. Further polarimetric observations are urged to monitor changes.”

Further photoelectric photometry (cf. IAUC 5818, 5829): June 15.232 UT, V = 9.94 ± 0.01, B–V = +0.38 ± 0.01 (D. H. Kaiser, Columbus, IN); July 7.98, V = 13.00 ± 0.05 (H. Mikuž, Ljubljana, Slovenia); 13.90, V = 13.19 ± 0.05 (Mikuž); 17.94, V = 13.95 ± 0.05 (Mikuž); 23.92, V = 14.13 ± 0.05 (Mikuž).

COMET MUELLER (1993a) Total magnitude estimates (cf. IAUC 5756): Apr. 8.86 UT, 13.8 (H.

Mikuž, Ljubljana, Slovenia, 0.20-m Baker-Schmidt camera + V filter + ST-6 CCD); 17.19, 13.1 (C. S. Morris, Pine Mountain Club, CA, 0.26m reflector, visual); May 22.85, 13.2 (Mikuž); July 29.08, 12.3 (Mikuž; sunward fan-like tail).

1993 August 4 Daniel W. E. Green